# Glimpse: Generalized Local Imaging with MLPs

### v2

@@ V Limitations and Conclusion
In the absence of calibration, we cannot expect to have paired ground truth images. In the following experiments, we only want to showcase the possibility to differentiably optimize over angles in Glimpse so we assume having access to paired data (while simulating the uncalibrated forward operator). In practice we could use a self-supervised loss, for example, based on equivariance [54].
@@
Figure 7: Estimated sensor geometry by Glimpse (LSG) and reconstructions for an uncalibrated system with a random sensor shift; as expected, the learnable sensor geometry can effectively learn the projection angles and exhibits excellent robustness with no degradation under such a big model mismatch and measurement noise (30dB).
@@
We assess the performance of Glimpse in situations with mismatched projection orientations. In the following experiments, we place $r=30$ sensors uniformly around the object at angles $\alpha=0^{\circ},6^{\circ},...,174^{\circ}$. We conduct a comparative analysis of three models: 1) Glimpse (vanilla), with no learnable sensor geometry, 2) Glimpse (LSG), incorporating the proposed learned sensor geometry, and 3) Glimpse (calibrated), operating under ideal conditions with no model mismatch (informed with correct projection angles). We let the Glimpse (LSG) learn the projection angles from the training data where the optimized values $\{\alpha_{m}\}_{i=1}^{r}$ obtained through training can provide a reliable estimate of the actual projection angles.
@@
Uncalibrated system with random sensor shifts: As shown in Figure 7(a), we randomly perturb projection angles by a normally distributed error so that $\alpha_{i}^{\text{given}}=\mathcal{N}(\alpha_{i},\sigma^{2})$; we set $\sigma=2^{\circ}$. We initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with $\alpha_{i}^{\text{given}}$. Figure 7(b) shows the estimated projection angles obtained through training—Glimpse (LSG) accurately recovers the angles even in the presence of 30 dB measurement noise. As shown in Figure 7(c), this accurate estimation of projection angles results in high-quality reconstructions by Glimpse (LSG) comparable with the network trained in an ideal calibrated system.
@@
Blind inversion with no information from projection angles: We consider the blind scenario where the model operates without any prior knowledge of the sensor geometry making inversion challenging. As shown in Figure 8(a), we initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with random values. The estimated projection angles are shown in Figure 8(b), highlighting Glimpse (LSG)’s ability for data-driven sensor geometry estimation. Figure 8(c) presents the reconstructions achieved by Glimpse in both its vanilla and LSG versions. As expected, FBP and the Glimpse (vanilla) show poor reconstructions due to the missing sensor geometry information. On the other hand, Glimpse (LSG) could accurately reconstruct both in-distribution and OOD samples. Remarkably, these results are comparable to those achieved by the calibrated Glimpse with informed projection angles.
@@
Figure 8: Estimated sensor geometry by Glimpse (LSG) and reconstructions for blind inversion; Glimpse (LSG) was initialized with random projection angles $\{\alpha_{m}\}_{i=1}^{r}$ (a) could reliably estimate the projection angles purely from data (b) resulting in high-quality reconstructions (c).
-C Optimal Filter for FBP Reconstruction

### v3

@@ 5.2 Looking forward: locality for other imaging modalities
In the absence of calibration, we cannot expect to have paired ground truth images. In the following experiments, we only want to showcase the possibility to differentiably optimize over angles in Glimpse so we assume having access to paired data (while simulating the uncalibrated forward operator). In practice, we could use a self-supervised loss, for example, based on equivariance [79].
Figure 12: The experimental arrangement for conducting uncalibrated imaging experiments
@@
Figure 13: Estimated sensor geometry by Glimpse (LSG) and reconstructions for an uncalibrated system with a random sensor shift; as expected, the learnable sensor geometry can effectively learn the projection angles and exhibits excellent robustness with no degradation under such a big model mismatch and measurement noise (30dB). We indicate PSNRs between the reconstructions and the ground truth.
@@
We assess the performance of Glimpse in situations with mismatched projection orientations. In the following experiments, we place $r=30$ sensors uniformly around the object at angles $\alpha=0^{\circ},6^{\circ},...,174^{\circ}$. We compare three models: 1) Glimpse (vanilla), with no learnable sensor geometry, 2) Glimpse (LSG), incorporating the proposed learned sensor geometry, and 3) Glimpse (calibrated), operating under ideal conditions with no model mismatch (informed with correct projection angles). Figure 12 demonstrates the experimental procedure for uncalibrated imaging experiments. We let the Glimpse (LSG) learn the projection angles from the training data where the optimized values $\{\alpha_{m}\}_{i=1}^{r}$ obtained through training can provide a reliable estimate of the actual projection angles.
.3.1 Uncalibrated system with random sensor shifts
@@
As shown in Figure 13(a), we randomly perturb projection angles by a normally distributed error so that $\alpha_{i}^{\text{given}}=\mathcal{N}(\alpha_{i},\sigma^{2})$; we set $\sigma=2^{\circ}$. We train Glimpse (vanilla) on this uncalibrated dataset; despite this mismatch in the forward operator, Glimpse (vanilla) can still generate high-quality reconstructions for in-distribution test data (only 0.6 dB drop compared to the calibrated system) as shown in the first row of the second column in Figure 13(c). However, the mismatch in the forward operator does not allow Glimpse (vanilla) to generalize well on OOD data (1.8 dB drop compared to the calibrated system) as shown in the second row of the second column in Figure 13(c). To address this issue, we initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with $\alpha_{i}^{\text{given}}$. Figure 13(b) shows the estimated projection angles obtained through training—Glimpse (LSG) accurately recovers the angles even in the presence of 30 dB measurement noise. As shown in Figure 13(c), this accurate estimation of projection angles results in high-quality reconstructions by Glimpse (LSG) comparable with the network trained in an ideal calibrated system.
.3.2 Blind inversion with no information from projection angles
@@
We consider the blind scenario where the model operates without any prior knowledge of the sensor geometry making inversion challenging. As shown in Figure 14(a), we initialize the projection angles $\{\alpha_{m}\}_{i=1}^{r}$ in the Glimpse (LSG) architecture with random values. The estimated projection angles are shown in Figure 14(b), highlighting Glimpse (LSG)’s ability for data-driven sensor geometry estimation. Figure 14(c) presents the reconstructions achieved by Glimpse in both its vanilla and LSG versions. As expected, FBP and the Glimpse (vanilla) show poor reconstructions due to the missing sensor geometry information. On the other hand, Glimpse (LSG) could accurately reconstruct both in-distribution and OOD samples. Remarkably, these results are comparable to those achieved by the calibrated Glimpse with informed projection angles.
@@
Figure 14: Estimated sensor geometry by Glimpse (LSG) and reconstructions for blind inversion; Glimpse (LSG) was initialized with random projection angles $\{\alpha_{m}\}_{i=1}^{r}$ (a) could reliably estimate the projection angles purely from data (b) resulting in high-quality reconstructions (c). We indicate PSNRs between the reconstructions and the ground truth.
.4 Optimal Filter for FBP Reconstruction